OpenStudy (anonymous):

sin^−1(sin(4pi/3))=??

OpenStudy (lgbasallote):

hint: \[\huge \sin^{-1} (\sin x) = x\] \[\huge \sin (\sin^{-1} x) = x\] does that help?

OpenStudy (anonymous):

nooo it doesnt equal x it equals -pi/3

OpenStudy (lgbasallote):

uhh you do realize what i wrote is different from what you wrote right?

OpenStudy (lgbasallote):

i was stating the "general rule"

OpenStudy (anonymous):

ok then it would be 4pi/3

jimthompson5910 (jim_thompson5910):

keep in mind that the range of arcsine is [-pi/2, pi/2]

jimthompson5910 (jim_thompson5910):

so -pi/3 is actually the answer

OpenStudy (lgbasallote):

if you simplify 4pi/3 you get -pi/3

jimthompson5910 (jim_thompson5910):

not quite, but sin(4pi/3) = sin(-pi/3)

OpenStudy (lgbasallote):

\[\huge \sin^{-1} (\sin (\frac{4\pi}{3}) \implies \sin^{-1} (\sin (-\frac{\pi}{3}))\] that'swhat i meant

OpenStudy (lgbasallote):

then following the rule i said... that becomes -pi3

OpenStudy (lgbasallote):

-pi/3*